To motorcycle traveling due north at 50 km/h the wind appears to come from northwest at 60 km/h. What is the true velocity of the wind
1
Expert's answer
2021-09-20T10:01:01-0400

The true velocity of the wind is


"\vec{w}=\frac{w_a-v}{\cos\theta}=\frac{60-50}{\cos45°}=14.1\text{ m/s W}"